OpenStudy (anonymous):

Shirt Color LargeMedium Total Red 42 48 90 Blue 35 40 75 Total 77 88 165 Which inference can be made from the data in the two-way table? P(shirt is red|shirt is large) = P(shirt is large) P(shirt is blue|shirt is large) = P(shirt is blue) P(shirt is medium and blue) = P(shirt is medium|shirt is blue) P(shirt is medium and blue) = P(shirt is medium) P(shirt is large|shirt is red) = P(shirt is red)

OpenStudy (anonymous):

I dont understand two way tables

OpenStudy (anonymous):

@NinjaDevo

OpenStudy (anonymous):

ehhh

OpenStudy (yanasidlinskiy):

Sorry...I was working on it a little. Size--large-medium--total Red - 42xxxx48xxxxx90 Blue--35xxxx 40xxxx 75 Total -77xxxx 88xxx 165 A) P( shirt is red / shirt is large) = 42 /77 P( shirt is large) = 77/ 165 A is false B) P(shirt is blue|shirt is large) = 35 / 77 P( shirt is blue) = 75/165 B is false C) P( shirt is medium and blue) = 40/165 P( shirt is medium / shirt is blue) = 40 / 75 C is false D) P( shirt is medium and blue) = 40/165 P( shirt is medium) = 88/165 E) P( shirt is large / shirt is red) = 42/90 P( shirt is red) = 90/ 165

OpenStudy (yanasidlinskiy):

\(\huge{\uparrow}\) That is what I came up with.

OpenStudy (anonymous):

So D or E ?

OpenStudy (yanasidlinskiy):

By the way... \(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\) And I'm going to go with D, because it doesn't really say what red is. Soo..yea.

OpenStudy (anonymous):

i think the answer is c P(shirt is medium and blue) = P(shirt is medium|shirt is blue)